OpenStudy (anonymous):

Please help !!!! 1. Randy drives his car to work each day. Suppose the expression (6/x) + (6/3x) represents the total time it takes Randy to drive to work in a day, where x represents the average speed of the car in miles per hour, during the first part of the drive. a. Use addition to simplify the expression. Show all the steps necessary to write the expression with a common denominator and to simplify the expression. b. If the speed of the car was 24 mph during the first part of the drive, find the total amount of time it took Randy to drive to work that day. please help

OpenStudy (solomonzelman):

you have to find the common denominator between these: \(\large\color{black}{ \displaystyle \frac{ 6}{x} +\frac{ 6}{3x} }\) so that you can add them.

OpenStudy (anonymous):

is it 3 ?